# Momentum and spin dynamics of photon in linear dielectric medium

### v1

@@ Abstract
Conclusion. In summary, we have analyzed the photon dynamics in a dielectric medium through the coupled Hamiltonian of photons and dipoles. The dipole oscillation in the medium mediates an indirect coupling between right-handed and left-handed photons, whose velocities are parallel and antiparallel to the spin $\mathbf{s}$, respectively.
@@
Conservation of spin ensures that the superposition of these photons reduces the photon velocity in the medium, with $s_{z}$ aligned parallel to the momentum operator $\hat{\mathbf{p}}$. The precession of transverse spin components $s_{x,y}$ around $\hat{\mathbf{p}}$ determines the photon phase. Refraction into the dielectric medium conserves the precession frequency and the tangential momentum $\hat{\mathbf{p}}_{t}$, thereby increasing the magnitude of $\hat{\mathbf{p}}$. This momentum corresponds to the Minkowski momentum, which sets the direction of spin and thus the direction of velocity.
Furthermore, the Abraham momentum governs the actual transfer of momentum to the medium, as the acceleration of the photon (Eq. 44) is directly related to the Lorentz force reaction on the dipole (Eq. 4). Our results clarify the coexistence of both Abraham and Minkowski momenta in dielectric media, assigning distinct physical roles to each.

### v3

@@ V Conclusion
In summary, we have analyzed the EM wave dynamics in a dielectric medium through the coupled Hamiltonian of EM waves and dipoles. The dipole oscillation in the medium mediates an indirect coupling between right-handed and left-handed EM waves, whose velocities are parallel and antiparallel to the spin $\mathbf{s}$, respectively. This helicity mixing is similar to the zitterbewegung effect known in relativistic quantum mechanics, where internal degrees of freedom generate rapid oscillatory motion. The single-particle picture of zitterbewegung reinforces the interpretation of light as possessing particle-like behavior, even in a classical-wave framework.
@@
Conservation of spin ensures that the superposition of these EM waves reduces the velocity of EM wave in the medium, with $s_{z}$ aligned parallel to the momentum operator $\hat{\mathbf{p}}$. The precession of transverse spin components $s_{x,y}$ around $\hat{\mathbf{p}}$ determines the EM wave phase. Refraction into the dielectric medium conserves the precession frequency and the tangential momentum $\hat{\mathbf{p}}_{t}$, thereby increasing the magnitude of $\hat{\mathbf{p}}$. This increase corresponds to the refractive index and gives rise to the Minkowski momentum, which determines the direction of propagation according to Snell’s law.
In contrast, the Abraham momentum represents the actual transfer of momentum to the medium. In our formulation, it is obtained as the expectation value of the spin-projected momentum vector and appears in the acceleration equation as the term balancing the Lorentz-force reaction on the dipoles. The Minkowski momentum, on the other hand, corresponds to the magnitude of the spin-projected momentum and governs canonical wave propagation. This coexistence mirrors the case of a relativistic Dirac electron, whose velocity magnitude equals the speed of light, while its expectation value is reduced by zitterbewegung. We also predict a zitterbewegung-like oscillation due to helicity mixing, which emerges from the dipole-mediated coupling of polarization modes. While this effect does not require a quantized field, it reflects internal dynamics typically associated with particle behavior, offering a conceptual bridge between classical wave motion and the particle picture. Our results clarify the coexistence of both Abraham and Minkowski momenta in dielectric media, assigning distinct physical roles to each.
Acknowledgments.
We acknowledge the support of Universitas Indonesia via PUTI Research Grant PKS-182/UN2.RST/HKP.05.00/2025